Solve graphically and check your answer algebraically.
X – Y = 2
2X + Y = 2
7X – 5Y = 6
1
Expert's answer
2020-08-04T19:23:06-0400


The system has no solution

"X=2+Y""2(2+Y)=2""7X\u22125Y=6"

"X=2+Y""Y=\u22121""7X\u22125Y=6"

"X=1""Y=-1""7X\u22125Y=6"

"X=1""Y=-1""7(1)\u22125(\u22121)=6,False"

Therefore the system has no solution.